find the difference quotient f(x+h)-f(x)/h for
f(x) = sqrt(x/(x+1)
please show work and explain this will be very helpful thank you so so much!

Answer by stanbon(75887)

find the difference quotient f(x+h)-f(x)/h for
f(x) = sqrt(x/(x+1)
------
f(x+h) = sqrt[(x+h)/(x+h+1)]
f(x) = sqrt(x/x+1)
------
Subtract to get: {sqrt[(x+h)/(x+h+1)] - sqrt[x/(x+1)]}/h
====================
Cheers,
Stan H.
=========